# Extended Emission Line Region in a Post-Starburst Galaxy Hosting TDE AT2019qiz and QPE

### v2

@@ 3.3 Gas and Stellar Kinematics
We measured the line-of-sight velocity and velocity dispersion of gas and stellar components using pPXF fitting. The emission lines H and H are associated with one kinematic component, [N ii], [S ii] and [O i] with another, and [O iii] as an independent component. The stellar continuum is treated as a separate kinematic component. The results are presented in Figure 6 for gas and Figure A.4 for stellar components as a comparison. In the gas velocity map, we overlay different line flux contours. The map reveals that the EELR region is nearly parallel to the kinematic major axis, seems to have non-circular motion compared to the stellar velocity map. Within the EELR region, the gas kinematic major axis is distorted, with the gas velocity map showing nearly zero line-of-sight velocity even outside the major axis. Perpendicular to the major axis, an additional kinematic component of gas is observed compared to the stellar velocity map, with one side moving towards us and the other moving away. The gas velocity exceeds the stellar component by approximately . The velocity dispersion map demonstrates lower gas dispersion within the EELR compared to its surroundings, contrasting with the centrally peaked stellar dispersion. Decoupling of gas and stellar motions has been observed in the EELR hosts of the TDE iPTF–16fnl (Wevers & French, 2024) and QPEs (Wevers et al., 2024), although our result is less pronounced. Several possibilities could explain these observations. First, the gas kinematics may be distorted by the galaxy bar potential (Sellwood & Sánchez, 2010; Holmes et al., 2015; Feng et al., 2022), while causing bar-induced inflow. Alternatively, the gas might originate from another galaxy through past minor merger events (Barnes & Hernquist, 1996; Barrera-Ballesteros et al., 2015), forming an misaligned gas disk (named “inner warp”). In this senario, the EELR gas resides above the galactic disk rather than within it, moving in a circular orbit with higher inclination than the stellar disk relative to the observer. This would cause the gas in the EELR region to primarily exhibit tangential velocity, resulting in nearly zero line-of-sight velocity, while being ionized by photons from the central AGN activity directed toward the observer. It is difficult to observationally distinguish between an inner warp and radial inflow (Zuleta et al., 2024), further detailed kinematic modeling is necessary to substantiate these hypotheses.
@@ 3.4 Ionizing Luminosity Estimation
Due to the finite speed of light, gas at different distances from the center is ionized by photons emitted during different epochs of AGN activity. Therefore, the EELR can reflect the history of AGN activity, or its duty cycle. By comparing the current AGN ionizing luminosity with the ionizing luminosity estimated from emission lines, we can preliminarily determine whether the emission lines can be ionized by the current AGN activity. Given that EUV is heavily absorbed by the interstellar medium, we use X-ray luminosity for an approximate estimation. Nicholl et al. (2020) observed with Swift XRT archival data that the luminosity in early 2019 was . In Nicholl et al. (2024), higher resolution Chandra observations revealed an additional X-ray source located southeast of the host galaxy nucleus. Excluding its influence, the quiescent X-ray luminosity of the nuclear region is . X-ray luminosity and UV ionizing luminosity are in different bands, therefore considering a bolometric correction factor of approximately 10 (Duras et al., 2020), the observed current quiescent bolometric luminosity is still lower than the ionizing luminosity inferred from the emission lines within the EELR (), but close to the ionizing luminosity estimated for the Seyfert region (). This suggests that the diffuse emission line region within the EELR cannot only be explained by the current nuclear activity, possibly indicating a decline in recent AGN activity, like recently faded AGN. However, the central region may be ionized by the remaining weak AGN activity, corresponding to a timescale of approximately 3000 years.
@@ 4 Summary
In this study, we present MUSE integral-field spectroscopy data analysis of an extended emission line region (EELR) in the host galaxy of TDE AT2019qiz followed by QPEs. The EELR extends to approximately 3.7 kpc, with a central compact region dominated by [O iii] emission and a diffuse [N ii] emission region in the outskirts.
@@
In TDE-QPE host galaxies, additional evidence for recent AGN activities has been inferred from QPE timing in the framework of EMRI+disk model (Zhou et al., 2024a, c, b). As in the unified scenario proposed by Jiang & Pan (2025), AGN activities may boost both the TDE rate and the formation rate of low-eccentricity EMRIs, consequently boost the occurrence of QPEs. Our finding of an EELR in the first clear TDE-QPE associated galaxy provides strong support for the unified model. In the near future, this scenario can be verified by more advanced IFS observations of the host galaxies of TDE-QPE associations, which are expected to be unveiled by various surveys in the golden era of time-domain astronomy. On the other hand, the detection of more distant EELRs requires excellent IFS capabilities with sufficient spatial resolution. Thus at present, only instruments like VLT/MUSE can effectively resolve these extended structures. Additionally, EELRs represent ideal targets for the upcoming China Space Station Telescope (CSST) (Zhan, 2011) integral-field spectrograph (CSST-IFS), which features a field of view, high spatial resolution of /spaxel, and wavelength coverage of 0.35-1.0 . This instrument will facilitate EELR observations of more TDE and QPE samples, enabling detailed emission line, stellar population, and kinematic measurements.

### v4

@@ 3.3 Gas and Stellar Kinematics
We measured the line-of-sight velocity and velocity dispersion of gas and stellar components using pPXF fitting. The emission lines H and H are associated with one kinematic component, [N ii], [S ii] and [O i] with another, and [O iii] as an independent component. The stellar continuum is treated as a separate kinematic component. The results are presented in Figure 6 for gas and Figure A.4 for stellar components as a comparison. In the gas velocity map, we overlay different line flux contours. The map reveals that the EELR region is nearly parallel to the kinematic major axis, seems to have non-circular motion compared to the stellar velocity map. Within the EELR region, the gas kinematic major axis is distorted, with the gas velocity map showing nearly zero line-of-sight velocity even outside the major axis. Perpendicular to the major axis, an additional kinematic component of gas is observed compared to the stellar velocity map, with one side moving towards us and the other moving away. The gas velocity exceeds the stellar component by approximately . The velocity dispersion map demonstrates lower gas dispersion within the EELR compared to the centrally peaked stellar dispersion. Decoupling of gas and stellar motions has been observed in the EELR hosts of the TDE iPTF–16fnl (Wevers & French, 2024) and QPEs (Wevers et al., 2024), although our result is less pronounced. Several possibilities could explain these observations. First, the gas kinematics may be distorted by the galaxy bar potential (Sellwood & Sánchez, 2010; Holmes et al., 2015; Feng et al., 2022), while causing bar-induced inflow. Alternatively, the gas might originate from another galaxy through past minor merger events (Barnes & Hernquist, 1996; Barrera-Ballesteros et al., 2015), forming a misaligned gas disk (named “inner warp”). In this scenario, the EELR gas resides above the galactic disk rather than within it, moving in a circular orbit with higher inclination than the stellar disk relative to the observer. This would cause the gas in the EELR region to primarily exhibit tangential velocity, resulting in nearly zero line-of-sight velocity, while being ionized by photons from the central AGN activity directed toward the observer. Additionally, since the observed emission lines are a brightness-weighted sum of ionized gas along the line of sight, geometric differences—such as a misalignment between the ionizing radiation cone and the stellar disk—can also affect the apparent velocity. It is difficult to observationally distinguish between an inner warp and radial inflow (Zuleta et al., 2024), further detailed kinematic modeling is necessary to substantiate these hypotheses.
@@ 3.4 Ionizing Luminosity Estimation
Due to the finite speed of light, gas at different distances from the center is ionized by photons emitted during different epochs of AGN activity. Therefore, the EELR can reflect the history of AGN activity, or its duty cycle. By comparing the current AGN ionizing luminosity with the ionizing luminosity estimated from emission lines, we can preliminarily determine whether the emission lines can be ionized by the current AGN activity. Given that EUV is heavily absorbed by the interstellar medium, we use X-ray luminosity for an approximate estimation. Nicholl et al. (2020) observed with the Swift X-ray Telescope (XRT) and reported that the luminosity in early 2019 was . In Nicholl et al. (2024), higher resolution Chandra observations revealed an additional X-ray source located southeast of the host galaxy nucleus. Excluding its influence, the quiescent X-ray luminosity of the nuclear region is . X-ray luminosity and UV ionizing luminosity are in different bands, therefore considering a bolometric correction factor of approximately 10 (Duras et al., 2020), the observed current quiescent bolometric luminosity is still lower than the ionizing luminosity inferred from the emission lines within the EELR (), but close to the ionizing luminosity estimated for the Seyfert region (). This suggests that the diffuse emission line region within the EELR cannot only be explained by the current nuclear activity, possibly indicating a decline in recent AGN activity, like recently faded AGN. However, the central region may be ionized by the remaining weak AGN activity, corresponding to a timescale of approximately 3000 years.
@@ 4 Summary
In this study, we present MUSE integral-field spectroscopy data analysis of an EELR in the host galaxy of TDE AT2019qiz followed by QPEs. The EELR extends to approximately 3.7 kpc, with a central compact region dominated by [O iii] emission and a diffuse [N ii] emission region in the outskirts.
@@
It should be noted that AGN as an ionizing source is merely an assumption here. Alternative sources such as past TDEs or evolved stars cannot be ruled out (as we discussed in Section 3.2), and the EELR may also result from a combination of multiple ionizing mechanisms. The primary aim of this study is to demonstrate the existence of the EELR in this galaxy, while additional observational evidence and more detailed modeling are needed to definitively determine the ionizing source. Nevertheless, considering the presence of mid-infrared echoes in the WISE light curve suggesting a dusty torus (Short et al., 2023; Pasham et al., 2025), the assumption of past AGN activity is reasonable.
In TDE-QPE host galaxies, additional evidence for recent AGN activities has been inferred from QPE timing in the framework of EMRI+disk model (Zhou et al., 2024a, b, 2025). As in the unified scenario proposed by Jiang & Pan (2025), AGN activities may boost both the TDE rate and the formation rate of low-eccentricity EMRIs, consequently boost the occurrence of QPEs. Our finding of an EELR in the first clear TDE-QPE associated galaxy provides strong support for the unified model. In the near future, this scenario can be verified by more advanced IFS observations of the host galaxies of TDE-QPE associations, which are expected to be unveiled by various surveys in the golden era of time-domain astronomy. On the other hand, the detection of more distant EELRs requires excellent IFS capabilities with sufficient spatial resolution. Thus at present, only instruments like VLT/MUSE can effectively resolve these extended structures. Additionally, EELRs represent ideal targets for the upcoming China Space Station Telescope (CSST) (Zhan, 2011; CSST Collaboration et al., 2025) integral-field spectrograph (CSST-IFS), which features a field of view, high spatial resolution of , and wavelength coverage of 0.35-1.0 . This instrument will facilitate EELR observations of more TDE and QPE samples, enabling detailed emission line, stellar population, and kinematic measurements.